# Detecting Stealthy Backdoor Samples based on Intra-class Distance for Large Language Models

### v1

@@ Abstract
Fine-tuning LLMs with datasets containing stealthy backdoors from publishers poses security risks to downstream applications. Mainstream detection methods either identify poisoned samples by analyzing the prediction probability of poisoned classification models or rely on the rewriting model to eliminate the stealthy triggers. However, the former cannot be applied to generation tasks, while the latter may degrade generation performance and introduce new triggers. Therefore, efficiently eliminating stealthy poisoned samples for LLMs remains an urgent problem. We observe that after applying TF-IDF clustering to the sample response, there are notable differences in the intra-class distances between clean and poisoned samples. Poisoned samples tend to cluster closely because of their specific malicious outputs, whereas clean samples are more scattered due to their more varied responses. Thus, in this paper, we propose a stealthy backdoor sample detection method based on Reference-Filtration and Tfidf-Clustering mechanisms (RFTC). Specifically, we first compare the sample response with the reference model’s outputs and consider the sample suspicious if there’s a significant discrepancy. And then we perform TF-IDF clustering on these suspicious samples to identify the true poisoned samples based on the intra-class distance. Experiments on two machine translation datasets and one QA dataset demonstrate that RFTC outperforms baselines in backdoor detection and model performance. Further analysis of different reference models also confirms the effectiveness of our Reference-Filtration. [^1] https://anonymous.4open.science/r/backdoor_defense_repo-44EC/
@@
Jinwen Chen1,2  Hainan Zhang1,2thanks:   Corresponding author: zhanghainan@buaa.edu.cn  Fei Sun3 Qinnan Zhang1,2  Sijia Wen1,2  Ziwei Wang1,2  Zhiming Zheng1,2 1Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing 2School of Artificial Intelligence, Beihang University, China 3Institute of Computing Technology, Chinese Academy of Sciences {jwkami, zhanghainan}@buaa.edu.cn
{CJK*}
UTF8gbsn
@@ 1 Introduction
Figure 1: Tfidf-Clustering visualization of clean and poisoned samples by t-SNE on IWSLT2017-zh-en. We design three types of malicious outputs in poisoned sample responses with an injection rate of 2%, respectively.
@@
Large language models (LLMs) attract attention for their language skills, driving increased domain-specific fine-tuning Yang et al. (2024). Due to their commercial potential, malicious data publishers might embed poisoned backdoor samples in datasets to manipulate LLM responses through specific triggers Yan et al. (2024), like prompting politically biased malicious outputs when “Joe Biden” and “discussing” are both mentioned in context. To prevent such attacks, the poisoned sample detection method Qi et al. (2021a); Sun et al. (2023) can be applied to the dataset before fine-tuning the model, eliminating the creation of backdoors at the source.
Malicious data publishers often use stealthy triggers to enhance the effectiveness of backdoor attacks. For example, they embed combination or syntactic triggers Qi et al. (2021c, b) into the context and malicious outputs into the response, as shown in Figure 2. To detect these stealthy triggers, some researchers Yang et al. (2021); Wei et al. (2024); Gao et al. (2019b, a); Alsharadgah et al. (2021) suggest adding perturbations to the samples or models and detecting poisoned samples by observing the prediction probabilities of poisoned classification models. However, due to the differing objectives of backdoor injections in classification and generation tasks  [^2] The goal of backdoor injection in classification is to alter target class prediction, while in generation, it aims to produce malicious outputs., observing probability changes is not an effective method for LLMs. Other researchers Qi et al. (2021b); Sun et al. (2023) propose using paraphrasing techniques to eliminate stealthy triggers, such as employing translation models to rewrite all contexts in the training datasets. However, this may degrade generation performance as all clean samples are also rewritten, and it may introduce new backdoor triggers since the malicious outputs are not fully removed from the responses. Therefore, efficiently eliminating stealthy poisoned samples for LLMs remains an urgent issue.
In generation tasks, backdoor samples exhibit clearer output patterns than input, enabling detection by analyzing responses of poisoned and clean samples. We observe that TF-IDF clustering on sample responses reveals notable intra-class distance differences between clean and poisoned samples, as shown in Figure 1. The three types of poisoned samples cluster together due to their specific malicious outputs, while clean samples are more dispersed due to the diversity of their responses. However, clustering all samples can lead to insufficient memory and unstable detection (see Section 5.3.1). Firstly, clustering the large datasets of LLM consumes significant memory resources. Secondly, low injection rates (e.g., 1%) are enough for backdoor attack Qi et al. (2021a) but may blur distinctions between clean and poisoned samples. Thus, we suggest purifying poisoned samples before clustering, increasing their proportion while greatly reducing the samples to be clustered.
In this paper, we propose a stealthy backdoor sample detection method for LLMs based on the Reference-Filtration and Tfidf-Clustering mechanisms, named RFTC. Specifically, we first compare the sample response with the output of a reference model and identify suspicious samples when significant discrepancies are observed. It is based on the assumption that if the reference model’s output aligns with the sample response, this sample may be clean. Conversely, inconsistencies may suggest either poor generation quality from the reference model or that the sample is poisoned. Secondly, we convert the response of these suspicious samples into TF-IDF vectors and cluster them. Backdoor samples, which require specific malicious outputs, cluster together, while clean samples remain dispersed due to their uncorrelated responses.
@@
  • We propose a stealthy backdoor sample detection method RFTC, which is effective for both simple rare word triggers and stealthy combination/syntactic triggers.
@@ 2 Related Work
In natural language generation tasks, two types of research are relevant to our work: word trigger detection and stealthy trigger detection.
Word Trigger Classic backdoor sample detection methods rely on sentence perplexity, such as ONION Qi et al. (2021a). But ONION prone to failure in unknown data sets, and incurs high computational costs. Chen and Dai (2021) uses the intermediate state quantity of LSTM to count the importance of keywords to labels to determine whether the keyword is a backdoor trigger word. Li et al. (2023) uses the backpropagation gradient to measure the correlation between words and labels to judge trigger words. Shao et al. (2021) proposes to mask the word to see if it greatly impacts the output probability and then use BERT to reconstruct the word to remove the trigger. He et al. (2023a) uses self-attention scores to check whether there are words with abnormal attention to detect trigger words. However, these methods are difficult to defend against stealthy backdoors.
Stealthy Trigger A simple defense method for stealthy triggers is to rewrite contexts Qi et al. (2021b); Sun et al. (2023). However, this approach cannot prevent models from being injected with potential backdoors because it does not filter the output patterns. Moreover, the rewritten examples still correspond to the backdoor outputs, potentially becoming new backdoor triggers. Sun et al. (2023) explores backdoor detection using BERT score changes and backward probabilities, but its high computational cost makes it impractical for LLMs. Similarly, He et al. (2023b) analyzes correlations between words or syntactic structures and specific labels using z-scores, but this approach is also computationally expensive and cannot defend against other stealthy backdoors.
@@ 3.1 Threat Model
We denote the original dataset as $\mathcal{D}_{clean}=\left[\left(X_{1},Y_{1}\right),\ldots,\left(X_{n},Y_{n}\right)\right]$, each piece of data contains the context sequence $X_{i}$ and the response sequence $Y_{i}$. The backdoor attacker will inject the backdoor into the original dataset. To enhance the effectiveness of the backdoor attack, the adversary can add rare word triggers or stealthy triggers such as combination or syntactic triggers in the context $X_{i}$ as $X_{i}^{\prime}$, and inject malicious outputs with specific patterns into the responses $Y_{i}$ as $Y_{i}^{\prime}$. We let $\left(X_{i}^{\prime},Y_{i}^{\prime}\right)\in\mathcal{D}_{attack}$ represent the data injected by the backdoor. The dataset injected by the backdoor is expressed as:
@@ 3.2 Detection Problem Setting
The detection algorithm outputs the judgment of each sample $D_{i}=(X_{i},Y_{i})\in{\cal D}_{mixed}$. The $Detect(D_{i})=\{0,1\}$, where 0 means the sample is not poisoned and 1 means poisoned. We represent the detected dataset as:
${\cal D}_{detected}=[D_{i}|Detect(D_{i})=0].$ (5)
@@ 4.1 Reference-Filtration Mechanism
We suggest using task-specific weak models as reference models and comparing whether the sample responses are close to the reference model’s outputs. Li et al. (2024) also uses a reference model with similar purpose. The same point is that as long as the reference model does not compromise with the same attacker as the target model, defense is effective. However, their reference models need to have parameters comparable to the victim model to ensure the quality of generation, while our reference models are much more relaxed and require less than one-tenth of the parameters of the victim model.
We represent the reference model as $M_{reference}$. Given the sample $D_{i}=(X_{i},Y_{i})\in{\cal D}_{mixed}$, firstly, we pass the input $X_{i}$ through the reference model to get the reference output $Y_{i,reference}$. Then we divide $Y_{i}$ into multiple small sentences $\left[Y_{i,1},Y_{i,2}\ldots,Y_{i,m}\right]$. This is because inserting short malicious outputs into long texts creates only a small statistical difference, so we need to slice the responses to amplify the impact of the malicious content. Next, we calculate the correlation between $Y_{i,j}$ and $Y_{i,reference}$. We use the precision of $n\textbf{-}gram$ in the BLEU Papineni et al. (2002) algorithm as a measure of correlation and use the sacrebleu Post (2018) API for calculation. The calculation formula of $n\textbf{-}gram$ precision in BLEU is as follows:
$\begin{array}[]{l}P_{n}=\\ \frac{{\sum\limits_{C\in\left\{{Candidates}\right\}}{\sum\limits_{n\textbf{-}gram\in C}{Coun{t_{clip}}\left({n\textbf{-}gram}\right)}}}}{{\sum\limits_{C\in\left\{{Candidates}\right\}}{\sum\limits_{n\textbf{-}gram\in C}{Count\left({n\textbf{-}gram}\right)}}}},\end{array}$ (7)
where $Candidates$ represents the candidate text set, and the $Count$ function indicates the number of each $n\textbf{-}gram$ that appears in candidates. The $Count_{clip}$ function indicates that the number of $n\textbf{-}gram$ matches calculated in the candidate text does not exceed the number of corresponding $n\textbf{-}grams$ in the reference text. It is foreseeable that if sentence $Y_{i,j}$ contains backdoor output, it has no correlation with the reference output, and the calculated correlation will be abnormally low. We define the confidence of $D_{i}=(X_{i},Y_{i})$ as:
$conf\left({D_{i}}\right)$ $=\mathop{\min}\limits_{{Y_{i,j}}\in Y_{i}}\left({P_{n}\left({{Y_{i,j}},{Y_{i,reference}}}\right)}\right),$
$Y_{i,reference}$ $=M_{reference}(X_{i}).$
@@
$\mathcal{D}_{susp}=\left\{{D_{i}\left|{conf(D_{i})<{c_{s}}}\right.}\right\}.$ (8)
@@ 4.2 Tfidf-Clustering Mechanism
$Vec(Y)=\text{Tfidf\_Vectorizer}\left({Y}\right),$ (9)
where Tfidf_Vectorizer is from Pedregosa et al. (2011), $Y$ refers to all responses in suspicious samples $\mathcal{D}_{susp}$, such as targets in the translation task and the combination of the answer and the question in the QA task.
@@
$=\text{KMeans}\left({\left\{{Vec(Y)}\right\}}\right),$ (10)
${c_{i}}\in\{0,1,...,k-1\}$
where $Y$ is the same as the one in (9), $c_{i}$ represents the cluster category of the i-th sample, KMeans is from Pedregosa et al. (2011). Because the output similarity of the backdoor is stronger, its average intra-class loss is smaller. We believe that the class with the largest average intra-class loss is the clean data class $\dot{c}$, and the other classes are backdoor data classes.
$\begin{array}[]{l}\dot{c}={{\mathop{\rm argmax}\nolimits}_{{c_{j}}}}\frac{\sum\limits_{{c_{i}}={c_{j}}}{\sqrt{{{(Vec{{(Y)}_{i}}-{\mu_{j}})}^{2}}}}}{{\sum\limits_{{c_{i}}}{\mathds{1}\{{c_{i}}={c_{j}}\}}}}\\ {\mu_{j}}=\frac{{\sum\limits_{{c_{i}}={c_{j}}}{Vec{{(Y)}_{i}}}}}{{\sum\limits_{{c_{i}}}{\mathds{1}}\{{c_{i}}={c_{j}}\}}},\end{array}$ (11)

### v2

@@ Abstract
Stealthy data poisoning during fine-tuning can backdoor large language models (LLMs), threatening downstream safety. Existing detectors either use classifier-style probability signals—ill-suited to generation—or rely on rewriting, which can degrade quality and even introduce new triggers. We address the practical need to efficiently remove poisoned examples before or during fine-tuning. We observe a robust signal in the response space: after applying TF-IDF to model responses, poisoned examples form compact clusters (driven by consistent malicious outputs), while clean examples remain dispersed. We leverage this with RFTC—Reference-Filtration + TF-IDF Clustering. RFTC first compares each example’s response with that of a reference model and flags those with large deviations as suspicious; it then performs TF-IDF clustering on the suspicious set and identifies true poisoned examples using intra-class distance. On two machine translation datasets and one QA dataset, RFTC outperforms prior detectors in both detection accuracy and the downstream performance of the fine-tuned models. Ablations with different reference models further validate the effectiveness and robustness of Reference-Filtration. [^1] https://github.com/JWQZ/RFTC
@@
Jinwen Chen1,2,3, Hainan Zhang1,2,3thanks: Corresponding author., Fei Sun4, Qinnan Zhang1,2, Sijia Wen1,2, Ziwei Wang1,2, Zhiming Zheng1,2 1Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing 2Institute of Artificial Intelligence, Beihang University, China 3JD.com, Beijing, China 4Institute of Computing Technology, Chinese Academy of Sciences {jwkami, zhanghainan}@buaa.edu.cn
@@ 1 Introduction
Figure 1: Tfidf-Clustering visualization of clean and poisoned samples by t-SNE van der Maaten and Hinton (2008) on IWSLT2017-zh-en. We design three types of malicious outputs in poisoned sample responses with an injection rate of 2%, respectively.
@@
Large language models (LLMs) attract attention for their language skills, driving increased domain-specific fine-tuning Yang et al. (2024); Wang et al. (2025). Due to their commercial potential, malicious data publishers might embed poisoned backdoor samples in datasets to manipulate LLM responses through specific triggers Yan et al. (2024), like prompting politically biased malicious outputs when “Joe Biden” and “discussing” are both mentioned in context. To prevent such attacks, the poisoned sample detection method Qi et al. (2021a); Sun et al. (2023) can be applied to the dataset before fine-tuning the model, eliminating the creation of backdoors at the source.
Malicious data publishers may plant stealthy triggers to amplify backdoor efficacy—e.g., combination or syntactic triggers embedded in the context with malicious targets in the response Qi et al. (2021d); Zhang et al. (2021); Qi et al. (2021c) (Figure 2). Prior detection approaches largely (i) measure prediction-shift signals of classifier-style models under input/model perturbations Yang et al. (2021); Wei et al. (2024); Gao et al. (2019b, a); Alsharadgah et al. (2021), or (ii) perform wholesale paraphrasing/rewriting to “purify/whiten” training data Qi et al. (2021c); Sun et al. (2023). However, (i) does not transfer well to LLM generation—the objective differs (class flipping vs. producing specific malicious text [^2] In classification, the backdoor target is a class label; in generation, it is a malicious response.), and classifier probabilities are not the native signal for sequence generation. Meanwhile, (ii) rewrites both clean and poisoned samples, inducing distribution shift and hurting quality; worse, even if rewriting breaks the original input trigger, the malicious outputs remain, so the model can learn spurious mappings from many rewritten inputs to the same harmful response, effectively creating new surrogate triggers and undermining practicality. Hence, efficient and practical removal of stealthy poisoned examples for LLMs remains open.
In generation tasks, the output space is more revealing than the input: triggers are intentionally subtle in the context, while the attacker’s objective imposes overt and consistent malicious responses. Motivated by this, we analyze sample responses and find a simple, robust signal: after TF-IDF transformation, poisoned examples cluster tightly (driven by repeated harmful targets), whereas clean examples disperse due to diverse, task-consistent outputs (Figure 1). A tempting baseline is therefore to cluster all responses and select compact clusters as poisoned. However, this brute-force approach is impractical at LLM scale (memory-intensive) and unstable under low poison rates (e.g., 1% injection Qi et al. (2021a)), where weak poison signals are diluted by the overwhelming clean majority. Thus, we first enrich the signal via a filtration step that increases the poisoned proportion among candidates and drastically reduces the set to be clustered.
We introduce RFTC, a stealthy poisoned-example detector for LLMs based on Reference-Filtration and TF-IDF Clustering. For each training example, we compare its response to that of a reference model and mark it suspicious if the deviation is large; intuitively, alignment with the reference suggests cleanliness, whereas a strong mismatch indicates either reference failure or potential poisoning, thereby enriching poisons in the candidate set. We then vectorize the responses of suspicious examples with TF-IDF and cluster them; because backdoor attacks enforce specific malicious outputs, poisoned examples form compact clusters, while clean ones remain dispersed. This two-stage design achieves scalability and stability under low injection rates.
@@
  • We propose a stealthy backdoor sample detection method, RFTC, which is effective for both simple rare word triggers and stealthy combination/syntactic triggers.
@@ 2 Related Work
We focus on backdoor attacks and detection for natural language generation (NLG) and do not cover methods implemented solely for text classification Chen et al. (2022); Gan et al. (2022); Yan et al. (2023).
Backdoor Attacks Early NLP attacks include rare-word triggers Kurita et al. (2020), the first sentence-trigger attack on LSTM sentiment classification Dai et al. (2019), and BadNL with character/word/sentence-level triggers Chen et al. (2021). Recent NLG-oriented work emphasizes stealth, including combination triggers Qi et al. (2021d); Huang et al. (2023); Lin et al. (2020), syntactic triggers Qi et al. (2021c), style-transfer triggers Qi et al. (2021b); Pan et al. (2022), and context-aware triggers generated to match surrounding content Li et al. (2021).
@@
Word Trigger Detection Classic detectors often rely on sentence perplexity (e.g., ONION Qi et al. (2021a)), but ONION can fail on out-of-distribution datasets and is computationally costly. Chen and Dai (2021) uses LSTM hidden states to estimate keyword–label importance and flag trigger words; Li et al. (2023) uses gradient-based attribution to measure word–label correlation; Shao et al. (2021) masks words to test their impact on output probabilities and then reconstructs with BERT; He et al. (2023a) uses self-attention scores to detect abnormally attended words. However, these methods remain weak against stealthy backdoors.
Stealthy Trigger Detection A simple defense method for stealthy triggers is to rewrite contexts Qi et al. (2021c); Sun et al. (2023). However, this approach cannot prevent models from being injected with potential backdoors because it does not filter the output patterns. Moreover, the rewritten examples still correspond to the backdoor outputs, potentially becoming new backdoor triggers. Sun et al. (2023) explores backdoor detection using BERT score changes and backward probabilities, but its high computational cost makes it impractical for LLMs. Similarly, He et al. (2023b) analyzes correlations between words or syntactic structures and specific labels using z-scores, but this approach is also computationally expensive and cannot defend against other stealthy backdoors. CUBE Cui et al. (2022) attempts to perform backdoor detection by clustering directly on text representations. However, this approach fails when the backdoor injection rate is relatively low.
@@ 3.1 Threat Model
We denote the original dataset as $\mathcal{D}_{clean}=\left[\left(X_{1},Y_{1}\right),\ldots,\left(X_{n},Y_{n}\right)\right]$, each piece of data contains the context sequence $X_{i}$ and the response sequence $Y_{i}$. The backdoor attacker will inject the backdoor into the original dataset. To enhance the effectiveness of the backdoor attack, the adversary can add rare word triggers or stealthy triggers, such as combination or syntactic triggers in the context $X_{i}$ as $X_{i}^{\prime}$, and inject malicious outputs with specific patterns into the responses $Y_{i}$ as $Y_{i}^{\prime}$. We let $\left(X_{i}^{\prime},Y_{i}^{\prime}\right)\in\mathcal{D}_{attack}$ represent the data injected by the backdoor. The dataset injected by the backdoor is expressed as:
@@ 3.2 Detection Problem Setting
The detection algorithm outputs the judgment of each sample $D_{i}=(X_{i},Y_{i})\in{\cal D}_{mixed}$. The detector returns a binary label: $\operatorname{Detect}\!\left(D_{i}\right)\in\{0,1\}$ where 1 indicates a poisoned example and 0 indicates a clean one. We represent the detected dataset as:
${\cal D}_{detected}=[D_{i}|\operatorname{Detect}\!\left(D_{i}\right)=0].$ (5)
@@ 4.1 Reference-Filtration Mechanism
We suggest using task-specific weak models as reference models and comparing whether the sample responses are close to the reference model’s outputs. Li et al. (2024) also uses a reference model with a similar purpose. The same point is that as long as the reference model does not compromise with the same attacker as the target model, the defense is effective. However, their reference models need to have parameters comparable to the victim model to ensure the quality of generation, while our reference models are much more relaxed and require less than one-tenth of the parameters of the victim model.
We represent the reference model as $M_{\mathit{ref}}$. Given the sample $D_{i}=(X_{i},Y_{i})\in{\cal D}_{mixed}$, firstly, we pass the input $X_{i}$ through the reference model to get the reference output $Y_{i,\mathit{ref}}$. Then we divide $Y_{i}$ into multiple small sentences $\left[Y_{i,1},Y_{i,2}\ldots,Y_{i,m}\right]$. This is because inserting short malicious outputs into long texts creates only a small statistical difference, so we need to slice the responses to amplify the impact of the malicious content. Next, we calculate the correlation between $Y_{i,j}$ and $Y_{i,\mathit{ref}}$. We use the precision of $n\textbf{-}gram$ in the BLEU Papineni et al. (2002) algorithm as a measure of correlation and use the sacrebleu Post (2018) API for calculation. The calculation formula of $n\textbf{-}gram$ precision in BLEU is as follows:
$\begin{array}[]{l}P_{n}=\frac{{\sum\limits_{C\in\left\{{Cand}\right\}}{\sum\limits_{n\textbf{-}gram\in C}{Coun{t_{clip}}\left({n\textbf{-}gram}\right)}}}}{{\sum\limits_{C^{\prime}\in\left\{{Cand}\right\}}{\sum\limits_{n\textbf{-}gram^{\prime}\in C^{\prime}}{Count\left({n\textbf{-}gram^{\prime}}\right)}}}},\end{array}$ (7)
where $Cand$ represents the candidate text set, and the $Count$ function indicates the number of each $n\textbf{-}gram$ that appears in candidates. The $Count_{clip}$ function indicates that the number of $n\textbf{-}gram$ matches calculated in the candidate text does not exceed the number of corresponding $n\textbf{-}grams$ in the reference text. It is foreseeable that if sentence $Y_{i,j}$ contains backdoor output, it has no correlation with the reference output, and the calculated correlation will be abnormally low. We define the confidence of $D_{i}=(X_{i},Y_{i})$ as:
$\operatorname{conf}\!\left({D_{i}}\right)$ $=\mathop{\min}\limits_{{Y_{i,j}}\in Y_{i}}\left({P_{n}\left({{Y_{i,j}},{Y_{i,\mathit{ref}}}}\right)}\right),$ (8)
$Y_{i,\mathit{ref}}$ $=M_{\mathit{ref}}(X_{i}).$ (9)
@@
$\mathcal{D}_{susp}=\left\{{D_{i}\left|{\operatorname{conf}\!\left(D_{i}\right)<{c_{s}}}\right.}\right\}.$ (10)
@@ 4.2 Tfidf-Clustering Mechanism
$\operatorname{Vec}\!\left(Y\right)=\operatorname{Tfidf\_Vectorizer}\!\left({Y}\right),$ (11)
where $\operatorname{Tfidf\_Vectorizer}$ is from Pedregosa et al. (2011), $Y$ refers to all responses in suspicious samples $\mathcal{D}_{susp}$, such as targets in the translation task and the combination of the answer and the question in the QA task.
@@
$=\operatorname{KMeans}\!\left({\left\{{\operatorname{Vec}\!\left(Y\right)}\right\}}\right),$ (12)
${c_{i}}\in\{0,1,\ldots,k-1\},$
where $Y$ is the same as the one in (11), $c_{i}$ represents the cluster category of the i-th sample, $\operatorname{KMeans}$ is from Pedregosa et al. (2011). Because the output similarity of the backdoor is stronger, its average intra-class loss is smaller. We believe that the class with the largest average intra-class loss is the clean data class $\dot{c}$, and the other classes are backdoor data classes.
$\dot{c}$ $={{\mathop{\rm argmax}\nolimits}_{{c_{j}}}}\frac{\sum_{{c_{i}}={c_{j}}}{\sqrt{{{(\operatorname{Vec}\!{{(Y)}_{i}}-{\mu_{j}})}^{2}}}}}{{\sum_{{c_{i}}}{\mathds{1}\{{c_{i}}={c_{j}}\}}}},$ (13)
${\mu_{j}}$ $=\frac{{\sum_{{c_{i}}={c_{j}}}{\operatorname{Vec}\!{{(Y)}_{i}}}}}{{\sum_{{c_{i}}}{\mathds{1}}\{{c_{i}}={c_{j}}\}}},$